OpenStudy (anonymous):

Simplify h(a+3)-h(3)/a when h(x)=x^2+4x+5

OpenStudy (anonymous):

I don't know what to do...

OpenStudy (psymon):

Ouch to that one. Well, you would start off with h(a+3). You need to plug (a+3) into every x you have in your function and simplify. If we can get that done okay we can keep going. But essentially, when it says f(__), whatever is in there is what gets replaced into every single variable in your equation.

OpenStudy (anonymous):

not nearly as bad as it looks what is \(f(3)\)?

OpenStudy (anonymous):

So I simplified it down to a^2+6x+4a/a but I think that it is wrong, because you cannot divide it by a still.

OpenStudy (anonymous):

or rather \(h(3)\) ?

OpenStudy (anonymous):

you have no \(x\) in this, so there is a mistake somewhere

OpenStudy (anonymous):

\(h(3)=9+12+5=26\) right?

OpenStudy (anonymous):

Wait, lol, I think that it is a^2+6a+4a/a

OpenStudy (anonymous):

well i don't know if that is right, but if it is, divide by \(a\) and get \(a+10\) lets try it

OpenStudy (anonymous):

\(h(3)=26\) \[h(a+3)=(a+3)^2+4(a+3)+5\\=a^2+6a+9+4a+12+5=a^2+10a+26\]

OpenStudy (anonymous):

so \(h(a+3)-h(3)=a^2+10a\)

OpenStudy (anonymous):

divide by each term by \(a\) and get \(a+10\)

OpenStudy (anonymous):

I see it... Lol I make the stupid mistakes always... That is what always gets me...

OpenStudy (anonymous):

i think you just forgot to combine like terms

OpenStudy (anonymous):

Nah, I'm just hardcore derping. It is the first day back to school, and the teacher gave us an assignment and said that there will be a quiz tomorrow, and I think that I am scared to take it tomorrow.